A gas at 25 °C fills a container whose volume is 1.05 x 103 cm3. The container plus gas has a mass of 837.6 g. The container, when emptied of all gas, has a mass of 836.2 g. What is the density of the gas at 25 °C? 


1
Expert's answer
2021-10-22T03:57:32-0400

Volume of Container = 1.05 * 103 cm3

Mass of empty Container = 836.2 g

Mass of container when gas is filled = 837.6 g


So, mass of gas = 837.6 - 836.2 = 1.4 g

Volume of gas = Volume of container = 1.05 * 103 cm3


Therefore, density of gas = Mass of gas / Volume of gas

Density = 1.4/1.05 * 103 = 1.3 * 10-3 g/cm3 = 0.0013 g/cm3


Hence, the density of gas is 0.0013 g/cm3.